OpenStudy (anonymous):

I need help!!!! solving ^3square root 2^3x =3/2

OpenStudy (anonymous):

I just want to make sure I'm reading this right... The problem is: [cube root of (2^(3x))] = 3/2 I'm not quite clear on what the problem is the way you wrote it...

OpenStudy (amistre64):

you should use paranethesis to indicate what is under the radical. 3rt(stuff under) is easier to understand

OpenStudy (anonymous):

cube root of 3^6 x^3 =3/2

OpenStudy (amistre64):

cbrt(3^6 * x^3) = 3/2 ; uncube the left side 3^2 * x = 3/2 ; divide out the 3^2 x = 3/2/(9) = 3/18 = 1/6 .. maybe

OpenStudy (anonymous):

First, start out by cubing both sides. This gives you 3^6 * x^3 = 27/8 Divide both sides by 3^6, which leaves: x^3 = 1 / (3^3 * 2^3) Then take the cube root of both sides for the answer: x = 1/(3*2) = 1/6.

OpenStudy (anonymous):

thank you